OpenStudy (tileer139):

Francine and Cheryl received equal scores on a test made up of multiple choice questions and an essay. Francine got 34 multiple choice questions correct and received 14 points for her essay. Cheryl got 30 multiple choice questions correct and received 22 points for her essay. How many points was each multiple choice question worth? points

OpenStudy (anonymous):

yes

OpenStudy (tileer139):

thanks

OpenStudy (anonymous):

anytime need anymore help?

OpenStudy (tileer139):

yeah can you

OpenStudy (tileer139):

i got 2 more

OpenStudy (anonymous):

go ahead

OpenStudy (tileer139):

In October, Greg and Thomas had the same amount of money in their savings accounts. In November, Greg deposited $120 into his account. Thomas increased the money in his account by 20%. When they compared their balances, they found that they were still equal. How much money did each person have in his account in October? $

OpenStudy (tileer139):

Two stores sell the same refrigerator for the same original price. Store A advertises that the refrigerator is on sale for 15% off the original price. Store B advertises that it is reducing the refrigerator’s price by $150. When Stephanie compares the sale prices of the refrigerator in both stores, she concludes that the sale prices are equal. Which equation models this situation? Let p represent the refrigerator’s original price.

OpenStudy (tileer139):

are u still there

OpenStudy (tileer139):

@Mehek14

OpenStudy (tileer139):

can u help

OpenStudy (anonymous):

yeah im here

OpenStudy (tileer139):

can u help

OpenStudy (anonymous):

yeah

OpenStudy (tileer139):

wats the answer

OpenStudy (anonymous):

hold on please

OpenStudy (tileer139):

ok

OpenStudy (anonymous):

do i get any medal for this?

OpenStudy (tileer139):

yeah

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

do u need anymore help in any subjects?

OpenStudy (tileer139):

hold on one sec

OpenStudy (anonymous):

okay

OpenStudy (tileer139):

Caleb and Sergio stacked boxes on a shelf. Caleb lifted 14 boxes and Sergio lifted 12 boxes. The boxes that Sergio lifted each weighed 10 lb more than the boxes Caleb lifted. Which expression represents the total number of pounds Sergio lifted? Let c represent the weight of each box Caleb lifted. A. 120c B. c + 10 C. 12(c + 10) D. 10(c − 12)

OpenStudy (tileer139):

i have 5

OpenStudy (anonymous):

okay hold on im fine with that i can help

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

next

OpenStudy (tileer139):

Two stores sell the same computer for the same original price. Store A advertises that the computer is on sale for 25% off the original price. Store B advertises that it is reducing the computer’s price by $180. When Brittany compares the sale prices of the computer in both stores, she concludes that the sale prices are equal. Which equation models this situation? Let p represent the computer’s original price. A. 0.25p = p – 180 B. 0.75p = p + 180 C. 0.25(p – 180) = p D. 0.75p = p − 180

OpenStudy (anonymous):

ok i got and web from oenstudy that i cant just give u the answer but what do u think it is

OpenStudy (tileer139):

b

OpenStudy (anonymous):

why do u think it's b